OpenStudy (anonymous):

Assume that the committee consists of 8 Republicans and 5 Democrats. A subcommittee of 4 is randomly selected from all subcommittees of 4 which contain at least 1 Democrat. What is the probability that the new subcommittee will contain at least 2 Democrats?

OpenStudy (anonymous):

good lord

OpenStudy (anonymous):

lets see if we can figure out how many subcommittees contain at least one democrat, and i guess the best way to do it is figure out how many contain no democrats and subtract it from the total possible subcommittees

OpenStudy (anonymous):

these things ar ekilling me and i have to have these done by 730 am

OpenStudy (anonymous):

number of subcommittees with not restriction is \[\dbinom{13}{4}=\frac{13\times 12\times 11\times 10}{4\times 3\times 2}=715\]

OpenStudy (anonymous):

number of subcommittees that are all republicans are \[\frac{8\times 7\times 6\times 5}{4!}=70\] so i guess we have \[715-70=645\] total subcommittees that contain at least one democrat

OpenStudy (anonymous):

that will be our denominator. now we have to figure out, of these 645 subcommittees with at least one democrat, how many have at least two

OpenStudy (anonymous):

ok exactly one democrat number of ways would be \[\dbinom{5}{1}\dbinom{8}{3}\] aka \[5\times 56=280\]

OpenStudy (anonymous):

so subtract this off from 645 to get the number that contain more than 1 democrat. you get \[645-280=365\] so our answer is \[\frac{365}{645}\]

OpenStudy (anonymous):

that didnt work

OpenStudy (anonymous):

you sure? i think the method is right. maybe a calculation error?

OpenStudy (anonymous):

the more i think about it the more i think it is right. do you have an answer?

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

whew!!

OpenStudy (anonymous):

im sorry not sure why it didnt work first time

OpenStudy (anonymous):

ok did you try the other one i sent? is this an on line class?

OpenStudy (anonymous):

no but home work is submitted online...

OpenStudy (anonymous):

oh i see. have fun!